# LLMs Think, But Not In Your Flow: Reasoning-Level Personalization for Black-Box Large Language Models

### v3

@@ Abstract
Large language models (LLMs) have recently achieved impressive performance across a wide range of natural language tasks and are now widely used in real-world applications. Among them, black-box LLMs—served via APIs without access to model internals—are especially dominant due to their scalability and ease of deployment. Despite their strong capabilities, these models typically produce generalized responses that overlook personal preferences and reasoning styles. This has led to growing interest in black-box LLM personalization, which aims to tailor model outputs to user-specific context without modifying model parameters. However, existing approaches primarily focus on response-level personalization, attempting to match final outputs without modeling personal thought process. To address this limitation, we propose RPM, a framework for reasoning-level personalization that aligns the model’s reasoning process with a user’s personalized logic. RPM first constructs statistical user-specific factors by extracting and grouping response-influential features from user history. It then builds personalized reasoning paths that reflect how these factors are used in context. In the inference stage, RPM retrieves reasoning-aligned examples for new queries via feature-level similarity and performs inference conditioned on the structured factors and retrieved reasoning paths, enabling the model to follow user-specific reasoning trajectories. This reasoning-level personalization enhances both predictive accuracy and interpretability by grounding model outputs in user-specific logic through structured information. Extensive experiments across diverse tasks show that RPM consistently outperforms response-level personalization methods, demonstrating the effectiveness of reasoning-level personalization in black-box LLMs.
@@ 1 Introduction
Recent advances in large language models (LLMs) have significantly improved performance across a wide range of natural language processing (NLP) tasks, from question-answering [1, 2] to recommendation [3, 4, 5, 6, 7, 8]. Most of these models deployed in practice are black-box models [9, 10, 11, 12], where access to internal parameters is unavailable. Despite their strong capabilities, these models typically produce generalized responses that do not reflect the goals, preferences, or reasoning styles of individual users. This limitation has led to growing interest in black-box LLM personalization [13, 14, 15, 16, 17], which aims to align model outputs with query and user-specific context without modifying the model parameters.
Recent approaches to black-box LLM personalization follow two main strategies: one retrieves personalized information, the other constructs tailored input prompts. The first approach selects information from user history, either randomly [4], via similarity-based retrieval [18], or by reranking based on utility scores [17]. The second approach injects user traits into the prompt through heuristic templates [18], profile summarization [8, 19, 20, 21, 22, 23] or iteratively optimize instruction [16]. However, both approaches share a fundamental limitation: they rely on unstructured raw data and align model outputs only at the response-level. They suffer from identifying which parts of the input contributed to the output, making it unclear whether the model’s response is grounded in the user’s personalized reasoning process, limiting both reliability and interpretability (Figure 1, Top).
@@
To address these limitations, our work aims to achieve reasoning-level personalization, improving both the quality of personalized responses and interpretability of black-box LLM personalization with given user history. Reasoning-level personalization refers to aligning the model’s reasoning process with the user’s personal reasoning process, in contrast to response-level alignment which focuses only on matching final responses without modeling how they were derived. Our key idea is to construct personalized reasoning from user history and guide inference on new queries by providing these prior reasoning paths (Figure 1, Bottom). By guiding black-box LLMs with reasoning paths constructed from user history, the model gains access to detailed references that reflect how similar situations were previously resolved based on each user’s reasoning logic. This allows the model to perform reasoning-aligned generation grounded in the user’s personalized reasoning for unseen queries.
To this end, we introduce RPM, Reasoning-Level Personalization for Black-Box LLM. To be specific, RPM performs reasoning-level personalization based on structured components and personalized reasoning derived from user history through three core stages. First, the model extracts response-influential features from the user history and groups them into user-specific meaningful statistical factors. Second, based on these components, it constructs personalized reasoning that aligns with the observed response from the user’s history. At inference time, the model generates output using feature-based retrieved examples, each paired with a personalized reasoning path. As shown in Figure 1, this structure not only enables the generation of personalized reasoning, but also provides effective criteria for retrieving useful reasoning examples. Furthermore, the retrieved reasoning paths allow the model to determine how to utilize structured components in unseen queries with personal logic, as they enable a reasoning process that clearly links the input context to the output.
We validate the effectiveness of RPM in personalizing black-box LLMs through extensive experiments across text classification, text generation, regression, and multiple-choice question-answering tasks. Compared to response-level alignment methods, RPM consistently yields higher personalization performance by modeling user-specific reasoning processes, demonstrating the effectiveness of reasoning-level personalization. In addition, human evaluation shows that structured components and personalized reasoning paths not only improve interpretability, but also increase users’ trust in the model’s outputs across diverse user scenarios. Additionally, experiments on example retrieval confirm that leveraging response-influential features enables more consistent identification of reasoning-aligned relevant examples, highlighting the practical utility of our retrieval mechanism.
Our main contributions are as follows: (1) We introduce personalized reasoning as a new paradigm for Black-Box LLM personalization, shifting the focus from response-level to reasoning-level personalization that captures how individuals think and respond. (2) We develop a novel framework that extracts response-influential features and user-specific statistical factors from the user history to effectively retrieve and guide personalized reasoning paths, enabling interpretable and consistent inference. (3) Through extensive experiments and human evaluation, we demonstrate that our approach improves personalization quality and interpretability across diverse tasks.
@@ 2.1 Retrieval-based LLM Personalization
A common strategy for black-box LLM personalization is to leverage a few past user information to condition model behavior during inference. In-Context Learning (ICL) achieves this by inserting several raw query–response pairs from a user’s history into the prompt [4, 3, 24, 25, 26], relying on the model’s in-context generalization ability to reproduce personalized responses. Retrieval-Augmented Generation (RAG) improves this approach by retrieving semantically similar examples from a longer history [18, 27, 28, 29], enabling scalability and better contextual fit. These approaches depend on query-level signals to retrieve user data, often including examples that appear relevant but offer little guidance on how to solve the current query. HYDRA [17] addresses this issue by training a reranker that reranks retrieved examples based on usefulness rather than similarity. However, the utility criteria require additional model parameter training, which increases complexity. In contrast, our method retrieves examples based on structured features that serve as the foundation for reasoning, offering an effective criteria for retrieveing useful reasoning exmples without additional model training.
@@ 2.2 Prompt-optimized LLM Personalization
Another approach to black-box LLM personalization involves refining prompts or profiles to better align model outputs with user preferences. Prompt engineering methods heuristically encodes user information into prompts [18, 30], while Fermi improves prompt quality via iterative updates based on model feedback [16]. Profile-based methods further summarize user traits into natural language descriptions appended to the prompt [19, 21, 23, 31]. However, even when effective prompts are found, these methods provide limited guidance on how to adaptively utilize input information in a case-specific way for each new query. Furthermore, it remains unclear how the provided input is utilized in generating the output, which lacks an explicit mapping between input attributes and reasoning steps. In contrast, our method conditions the model on examples containing user-specific reasoning paths constructed from user history, enabling reasoning-level personalization with personal logic. Through structured features and factors, our approach produces interpretable responses that reveal how specific query components contribute to the model’s predictions.
@@ 3 RPM: Reasoning-Level Personalization for Black-Box LLMs
In this section, we present RPM, a framework for reasoning-level personalization of black-box LLMs using structured user information. RPM aligns model inference with user-specific reasoning logic by constructing and leveraging personalized reasoning paths from history. The framework consists of three key components: (1) personalized factor construction, which extracts and groups response-relevant features into statistical user-level factors (Section 3.2); (2) personalized reasoning construction, which builds personalized reasoning paths based on user’s past responses (Section 3.3); and (3) reasoning-aligned generation, which retrieves and applies these reasoning paths to generate accurate and interpretable outputs (Section 3.4). Figure 2 shows the overview of our RPM framework.
@@ 3.2 Personalized Factor Construction
Structured feature extraction. Given a user’s history $\mathcal{H}_{u}$, we extract features from each query $q_{i}$ to identify elements that may influence the user’s response $a_{i}$. A raw query may contain elements that are irrelevant to the prediction, introducing noise into LLM personalization. To address this, we utilize $\mathcal{M}$ to extract features from each query $q_{i}$ that are likely to influence the user’s response, leveraging the LLM’s ability to capture semantically rich and relevant signals from text [32, 33]. Specifically, we prompt the LLM with the feature extraction instruction for the query $q_{i}$, resulting in a set of features $\mathcal{G}_{q_{i}}=\{f_{j}\}_{j=1}^{|\mathcal{G}_{q_{i}}|}$ defined as follows:
$\mathcal{G}_{q_{i}}=\mathcal{M}(q_{i}),\quad f_{j}=(\texttt{name}_{j},\texttt{context}_{j},\texttt{factor}_{j}),$ (1)
where $f_{j}$ denotes a $j$-th feature extracted from $q_{i}$, with namej representing the semantic label of the feature, contextj specifying the context that clarifies the feature’s intended meaning, and factorj denotes a placeholder indicating the latent factor to which the feature is likely to belong. This process yields a query-specific feature set $\mathcal{G}_{q_{i}}$ by identifying potentially influential elements from $q_{i}$ that are explicitly recognizable to the language model. These extracted features not only reduce the noise and ambiguity of the raw query, but also serve as foundations for supporting both downstream factor construction and reasoning over user’s underlying patterns. Additional experiments on feature composition are provided in the Appendix D.2.
Factor generation via LLM-based clustering. To capture generalized user-level behavior patterns, we group the extracted features across all queries $\{q_{i}\}_{i=1}^{N}$ into a set of factors, which are higher-level semantic clusters. We adopt the LLM-based clustering method proposed in [34, 35, 36, 37], which groups features based on goal-related semantic similarity. This method aligns well with our objective, as semantically coherent clusters often reflect consistent reasoning tendencies unique to each user. Formally, the clustering process produces a set of user-specific factors:
$\{F^{(m)}\}_{m=1}^{M}=\texttt{LLM\_Cluster}\left(\bigcup_{i=1}^{N}\mathcal{G}_{q_{i}}\right).$ (2)
Here, each factor $F^{(m)}=\{f_{j}\}_{j=1}^{|F^{(m)}|}$ represents a semantically coherent subset of features identified by the clustering algorithm. This LLM-driven semantic clustering transforms low-level, query-specific features into structured, user-level representations that not only capture recurring semantic patterns but also serve as containers for storing quantitatively analyzable statistics. We include more detailed explanation about clustering in Appendix B.
@@
The resulting metric values computed in Equations (3)–(6) are stored as $\theta^{(m)}$. Together, these statistics provide a quantifiable profile for each factor, characterizing how frequently it appears, how often it meaningfully contributes to predictions, and in what direction it tends to influence the response. Please refer to Appendix F and G for the constructed features, factors and prompts.
@@ 3.3 Personalized Reasoning Construction
Once features and factors are extracted from user history, we generate a personalized reasoning path for each query-response pair to capture how the given information can be utilized to produce the correct response in an interpretable manner. While the extracted features and aggregated factors provide a rich representation of user-specific information, they do not explicitly indicate how these elements influence the final prediction. To address this, we prompt an LLM $\mathcal{M}$ with the reasoning instruction, query $q_{i}$, its associated features $\mathcal{G}_{q_{i}}$, user-level factors $\mathcal{C}_{u}$, and the corresponding response $a_{i}$, and instruct it to generate a reasoning path that connects the relevant information to the response: $r_{q_{i}}=\mathcal{M}(q_{i},\mathcal{G}_{q_{i}},\mathcal{C}_{u},a_{i})$ where $r_{q_{i}}$ denotes the generated reasoning path that explains the user’s reasoning logic based on provided query elements. Each reasoning path is then stored in the user’s memory $\mathcal{S}_{u}$, referred to as reasoning-augmented user history, as part of a tuple containing all relevant components: $\mathcal{S}_{u}=\left\{(q_{i},\mathcal{G}_{q_{i}},r_{q_{i}},a_{i})|(q_{i},a_{i})\in\mathcal{H}_{u}\right\}.$ This reasoning augmentation allows the framework to retain query-level reasoning behavior, enabling inference to be guided by personalized examples that encapsulate both structural information and user-aligned reasoning logics. Please refer to Appendix F and G for generated reasonings and prompts.
@@ 3.4 Reasoning-Aligned Generation
Factor-aware feature extraction. Following the procedure described in Section 3.2, we extract features from the target query $q\textquoteright$ ’ with reference to the user-specific factor set $\mathcal{C}_{u}$. Each extracted feature contains an associated factor field, which allows the model to leverage the corresponding statistical summaries during generation, enabling reasoning grounded in personal behavior patterns.
Retrieving useful reasoning examples. We retrieve reasoning examples from the augmented user history $\mathcal{S}_{u}$ that are useful for reasoning-level personalization. Our retrieval process formulates retrieval criterion by using $\mathcal{G}_{q^{\prime}}$ that includes response-influential elements. This feature-based formulation provides an effective criterion for selecting personalized reasoning examples. Specifically, we compute the semantic similarity between $\mathcal{G}_{q^{\prime}}$ and each stored $\mathcal{G}_{q}$, and retrieve the top-$K$ most relevant examples:
$\mathcal{S}_{u,q^{\prime}}^{\texttt{ret}}=\{(q,\mathcal{G}_{q},r_{q},a)\in\mathcal{S}_{u}|\text{Top-K}\,\cos\left(f(\mathcal{G}_{q^{\prime}}),f(\mathcal{G}_{q})\right)\},$ (7)
where $f(\mathcal{G}_{q})$ is the embedding of the concatenated feature texts and $\cos(\cdot)$ is cosine similarity. Each retrieved example $(q,\mathcal{G}_{q},r_{q},a)$ provides structured logics on how similar user-specific information was processed in the past.
Reasoning example-augmented generation. We then guide the black-box LLM $\mathcal{M}$ using the retrieved examples. Standard few-shot prompting often provides relevant examples but lacks guidance on how the given information should be interpreted to reach the correct response. In contrast, our approach supplies not only structured features and factors but also personalized reasoning paths, illustrating how such information has been used in prior response generation. Formally, we perform inference by promtping $\mathcal{M}$ on the target query $q^{\prime}$, its extracted representations $\mathcal{G}_{q^{\prime}}$, user-specific factor $\mathcal{C}_{u}$, and the retrieved reasoning examples $\mathcal{S}_{u,q^{\prime}}^{\texttt{ret}}$, i.e., $r_{q^{\prime}},a^{\prime}=\mathcal{M}(q^{\prime},\mathcal{G}_{q^{\prime}},\mathcal{C}_{u},\mathcal{S}_{u,q^{\prime}}^{\texttt{ret}}).$ By incorporating the personalized reasoning examples, the model is better able to interpret user-specific reasoning logic, enabling both effective personalization and transparent interpretation grounded in structured components and previous reasoning paths. Please refer to Appendix F and G for examples and prompts.

### v4

# RPM: Reasoning-Level Personalization for Black-Box Large Language Models

@@ Abstract
While black-box large language models are widely deployed, they produce generic outputs that overlook individual user preferences. Current personalization methods are fundamentally limited to response-level personalization; they only match final outputs, failing to model the underlying reasoning that connects user behavior to responses. To address this, this work introduces reasoning-level personalization as a new paradigm and proposes RPM, the first systematic framework designed to guide the model’s reasoning process using structured rationales constructed from patterns in a user’s behavior. RPM constructs a structured model of user behavior—built from response-influential features and statistical factors—to create personalized reasoning paths and retrieve beneficial examples for guiding inference through a feature-based retrieval mechanism. Extensive experiments across four diverse tasks demonstrate that RPM consistently outperforms existing response-level methods while simultaneously enhancing both personalization performance and interpretability, providing a promising direction for black-box LLM personalization.
@@ 1 Introduction
Recent advances in large language models (LLMs) have significantly improved performance across a wide range of natural language processing tasks (Hendrycks et al., 2020; Liu et al., 2023; Dai et al., 2023; Liu et al., 2024). Most deployed models operate as black-box systems where internal parameters are inaccessible (Brown et al., 2020; Achiam et al., 2023; Team et al., 2024), posing a fundamental challenge to providing personalized responses that align with individual user preferences and behavioral patterns. This has led to growing interest in black-box LLM personalization (Zhang et al., 2024; Kirk et al., 2023; Kim & Yang, 2024; Zhuang et al., 2024), which aims to tailor model outputs to user-specific contexts without modifying the model parameters.
Current black-box LLM personalization approaches fall into two main categories. Retrieval-based methods, which select historical data via similarity (Salemi et al., 2023) or utility scoring (Zhuang et al., 2024), and prompting-based methods, which refine inputs through heuristic templates (Salemi et al., 2023) or iterative updates (Kim & Yang, 2024). However, both strategies share a fundamental limitation: they focus exclusively on Response-Level Personalization, where the objective is limited only to matching the final output (Figure 1, Top). This limitation creates two key challenges. First, Superficial Pattern Learning. These systems can only learn shallow correlations between the overall input and the final output, failing to capture how specific components within the input influence the response. Second, Lack of Interpretability. Without an explicit reasoning path, it is hard to determine whether the model’s output reflects authentic user preferences or misleading correlations, which threatens the reliability of the system.
@@
In response to these limitations, we propose Reasoning-Level Personalization as a new paradigm that aims to model the reasoning process inferred from user behaviors. While this paradigm holds the potential for deeper behavioral insight and interpretability, realizing these advantages is a significant challenge. Our experiments confirm this: the most straightforward approach, applying zero-shot chain-of-thought (CoT) prompting (Kojima et al., 2022), yields inconsistent performance compared to zero-shot baselines. A more advanced approach, constructing reasoning paths for relevant historical data as few-shot CoT examples, also fails to improve upon few-shot baselines. These failures highlight a critical gap: the absence of a systematic framework capable of transforming raw behavioral patterns into a structured reasoning model that the LLM can reliably follow.
To address this critical need, this paper introduces RPM, Reasoning-Level Personalization for Black-Box LLM, a novel framework that aligns an LLM’s reasoning process with a user’s behavior by constructing and leveraging structured, personalized reasoning paths derived from their history (Figure 1, Bottom). This is achieved through four key innovations: (1) Instead of naively summarizing user history into a profile (Richardson et al., 2023; Kim et al., 2024b; Sun et al., 2024; Kim et al., 2025b), the framework constructs a structured user model by extracting response-influential features from each interaction, grouping them into quantifiable factors, and assigning statistical meaning. (2) It then builds personalized reasoning based on this structured model. Providing these reasoning-augmented examples during inference effectively aligns the LLM’s logic with the user’s, leading to significant performance gains. (3) This strong alignment is further enhanced by feature-based retrieval, a mechanism that effectively retrieves beneficial samples based on features, providing a more relevant foundation for reasoning than simple topical matching. (4) Finally, the framework achieves high interpretability by grounding its outputs in structured reasoning paths, making individual components like features and factors explicit in the final reasoning process.
In summary, our contributions are threefold: (1) The introduction and formalization of reasoning-level personalization, a new paradigm that shifts the focus from matching final responses to aligning the underlying reasoning process of LLMs with a user’s specific behavior pattern. (2) The proposal of RPM, a novel and systematic framework that effectively guides an LLM’s inference with high interpretability in black-box LLM personalization. (3) Comprehensive empirical evidence validating the framework’s effectiveness. The experiments not only demonstrate its state-of-the-art performance but also confirm the significant contribution of its core technical components. Furthermore, human evaluations verify that RPM ’s outputs are significantly more interpretable and trustworthy.
@@ 2.1 Retrieval-based LLM Personalization
A common strategy for black-box LLM personalization is to leverage a few past user information to condition model behavior during inference. In-Context Learning (ICL) achieves this by inserting several raw query–response pairs from a user’s history into the prompt (Dai et al., 2023; Liu et al., 2023; Kang et al., 2023; Di Palma et al., 2023; Wang & Lim, 2023), relying on the model’s in-context generalization ability to reproduce personalized responses. Retrieval-Augmented Generation (RAG) improves this approach by retrieving semantically similar examples from a longer history (Salemi et al., 2023; 2024; Gao et al., 2023; Li et al., 2023), enabling scalability and better contextual fit. These approaches depend on query-level signals to retrieve user data, often including examples that appear relevant but offer little guidance on how to solve the current query. HYDRA (Zhuang et al., 2024) addresses this issue by training a reranker that reranks the retrieved examples based on usefulness rather than similarity. However, the utility criteria require additional model parameter training, which increases complexity. In contrast, our method retrieves examples based on structured features that serve as the foundation for reasoning, offering an effective criterion for retrieving useful reasoning examples without requiring additional model training.
@@ 2.2 Prompt-optimized LLM Personalization
Another approach to black-box LLM personalization involves refining prompts or profiles to better align model outputs with user preferences. Prompt engineering methods heuristically encode user information into prompts (Salemi et al., 2023; Hwang et al., 2023), while Fermi improves prompt quality via iterative updates based on model feedback (Kim & Yang, 2024). Profile-based methods further summarize user traits into natural language descriptions appended to the prompt (Richardson et al., 2023; Kim et al., 2024b; 2025b; ). However, even when effective prompts are found, these methods provide limited guidance on how to adaptively utilize input information in a case-specific way for each new query. Furthermore, it remains unclear how the provided input is utilized in generating the output, which lacks an explicit mapping between input attributes and reasoning steps. In contrast, our method conditions the model on examples containing user-specific reasoning paths constructed from user history, enabling reasoning-level personalization with personal behavior pattern. Through structured features and factors, our approach produces interpretable responses that reveal how specific query components contribute to the model’s predictions.
@@ 3 RPM: Reasoning-Level Personalization for Black-Box LLMs
In this section, we present RPM, a framework for reasoning-level personalization of black-box LLMs using structured user information. RPM aligns model inference with user-specific behavior pattern by constructing and leveraging personalized reasoning paths from history. The framework consists of three key components: (1) personalized factor construction, which extracts and groups response-relevant features into statistical user-level factors (Section 3.2); (2) personalized reasoning construction, which builds personalized reasoning paths based on user’s past responses (Section 3.3); and (3) reasoning-aligned generation, which retrieves and applies these reasoning paths to generate accurate and interpretable outputs (Section 3.4). Figure 2 shows the overview of our RPM framework. For details on the notation and algorithms, please refer to Appendix A and Appendix B.
@@ 3.2 Personalized Factor Construction
Structured feature extraction. Given a user’s history $\mathcal{H}_{u}$, we extract features from each query $q_{i}$ to identify elements that may influence the user’s response $a_{i}$. A raw query may contain elements that are irrelevant to the prediction, introducing noise into LLM personalization. To address this, we utilize $\mathcal{M}$ to extract features from each query $q_{i}$ that are likely to influence the user’s response, leveraging the LLM’s ability to capture semantically relevant signals from text (Kim et al., 2024a; Seo et al., 2024). Specifically, we prompt the LLM with the feature extraction instruction for the query $q_{i}$, resulting in a set of features $\mathcal{G}_{q_{i}}=\{f_{j}\}_{j=1}^{|\mathcal{G}_{q_{i}}|}$ defined as follows: $\mathcal{G}_{q_{i}}=\mathcal{M}(q_{i}),\quad f_{j}=(\texttt{name}_{j},\texttt{context}_{j},\texttt{factor}_{j})$, where $f_{j}$ denotes a $j$-th feature extracted from $q_{i}$, with namej representing the semantic label of the feature, contextj specifying the context that clarifies the feature’s intended meaning, and factorj denotes a placeholder indicating the latent factor to which the feature is likely to belong. This process yields a query-specific feature set $\mathcal{G}_{q_{i}}$ by identifying potentially influential elements from $q_{i}$ that are explicitly recognizable to the language model. These extracted features not only reduce the noise and ambiguity of the raw query, but also serve as foundations for supporting both downstream factor construction and reasoning over user’s underlying patterns. Additional experiments on feature composition are provided in the Appendix E.2.
Factor generation via LLM-based clustering. To capture generalized user-level behavior patterns, we group the extracted features across all queries $\{q_{i}\}_{i=1}^{N}$ into a set of factors, which are higher-level semantic clusters. We adopt the LLM-based clustering method proposed in Wang et al. (2023), which groups features based on goal-related semantic similarity. This method aligns well with our objective, as semantically coherent clusters often reflect consistent reasoning tendencies unique to each user. Formally, the clustering process produces a set of user-specific factors: $\{F^{(m)}\}_{m=1}^{M}=\texttt{LLM\_Cluster}\left(\bigcup_{i=1}^{N}\mathcal{G}_{q_{i}}\right).$ Here, each factor $F^{(m)}=\{f_{j}\}_{j=1}^{|F^{(m)}|}$ represents a semantically coherent subset of features identified by the clustering algorithm. This LLM-driven semantic clustering transforms low-level, query-specific features into structured, user-level representations that not only capture recurring semantic patterns but also serve as containers for storing quantitatively analyzable statistics. More detailed explanation about clustering is in Appendix C.
@@
The resulting metric values computed in Equations (1)–(4) are stored as $\theta^{(m)}$. Together, these statistics provide a quantifiable profile for each factor, characterizing how frequently it appears, how often it meaningfully contributes to predictions, and in what direction it tends to influence the response. The validity of the extracted features and aggregated factors is evaluated in Section 4.4.
@@ 3.3 Personalized Reasoning Construction
Once features and factors are extracted from user history, we generate a personalized inference path for each query-response pair to capture how the given information leads to the observed response in an interpretable manner. While the extracted features and aggregated factors provide a rich representation of user-specific information, they do not explicitly indicate how these elements influence the final prediction. To address this, we prompt an LLM $\mathcal{M}$ with the reasoning instruction, query $q_{i}$, its associated features $\mathcal{G}_{q_{i}}$, user-level factors $\mathcal{C}_{u}$, and the corresponding response $a_{i}$, and instruct it to generate a reasoning path that connects the relevant information to the response: $r_{q_{i}}=\mathcal{M}(q_{i},\mathcal{G}_{q_{i}},\mathcal{C}_{u},a_{i})$ where $r_{q_{i}}$ denotes the generated reasoning path that explains the user’s behavioral pattern based on provided query elements. Each reasoning path is then stored in the user’s memory $\mathcal{S}_{u}$, referred to as reasoning-augmented user history, as part of a tuple containing all relevant components: $\mathcal{S}_{u}=\left\{(q_{i},\mathcal{G}_{q_{i}},r_{q_{i}},a_{i})|(q_{i},a_{i})\in\mathcal{H}_{u}\right\}.$ This reasoning augmentation allows the framework to retain query-level reasoning behavior, enabling inference to be guided by personalized examples that encapsulate structural information and user-aligned behavior patterns.
@@ 3.4 Reasoning-Aligned Generation
Factor-aware feature extraction. Following the procedure described in Section 3.2, we extract features from the target query $q$’ with reference to the user-specific factor set $\mathcal{C}_{u}$. Each extracted feature contains an associated factor field, which allows the model to leverage the corresponding statistical summaries during generation, enabling reasoning grounded in personal behavior patterns.
Retrieving useful reasoning examples. We retrieve reasoning examples from $\mathcal{S}_{u}$ that are useful for reasoning-level personalization. Our retrieval process formulates retrieval criterion by using $\mathcal{G}_{q^{\prime}}$ that includes response-influential elements. This feature-based formulation provides an effective criterion for selecting personalized reasoning examples. Specifically, we compute the semantic similarity between $\mathcal{G}_{q^{\prime}}$ and each stored $\mathcal{G}_{q}$, and retrieve the top-$K$ most relevant examples: $\mathcal{S}_{u,q^{\prime}}^{\texttt{ret}}=\{(q,\mathcal{G}_{q},r_{q},a)\in\mathcal{S}_{u}|\text{Top-K}\,\cos\left(f(\mathcal{G}_{q^{\prime}}),f(\mathcal{G}_{q})\right)\}(5)$, where $f(\mathcal{G}_{q})$ is the embedding of the concatenated feature texts and $\cos(\cdot)$ is cosine similarity. Each retrieved example $(q,\mathcal{G}_{q},r_{q},a)$ provides structured logics on how similar user-specific information was processed in the past.
Reasoning example-augmented generation. We then guide the black-box LLM $\mathcal{M}$ using the retrieved examples. Standard few-shot prompting often provides relevant examples but lacks guidance on how the given information should be interpreted to reach the correct response. In contrast, our approach supplies not only structured features and factors but also personalized reasoning paths, illustrating how such information has been used in prior response generation. Formally, we perform inference by promtping $\mathcal{M}$ on the target query $q^{\prime}$, its extracted representations $\mathcal{G}_{q^{\prime}}$, user-specific factor $\mathcal{C}_{u}$, and the retrieved reasoning examples $\mathcal{S}_{u,q^{\prime}}^{\texttt{ret}}$, i.e., $r_{q^{\prime}},a^{\prime}=\mathcal{M}(q^{\prime},\mathcal{G}_{q^{\prime}},\mathcal{C}_{u},\mathcal{S}_{u,q^{\prime}}^{\texttt{ret}}).$ By incorporating the personalized reasoning examples, the model is better able to interpret user-specific behavior pattern, enabling both effective personalization and transparent interpretation grounded in structured components and previous reasoning paths.
Prompt Design Our prompts are based on a generalizable template. The core structure that guides the model’s reasoning remains consistent across all tasks. Task-specificity is confined to designated placeholders for task descriptions, ensuring the framework is broadly applicable with minimal modification, contrary to the concern that they are heavily tailored. Therefore, our template-based approach provides a principled and scalable strategy, enabling the framework to be generalized to new tasks with minimal, targeted adjustments. Please refer to Appendix G and H for examples.